# Consensus formation in quality-sensitive interdependent agent systems

### v1

@@ III.2 Asymmetric discovery scenario
As before, we can quantify the outcome of the decision process in terms of a consensus parameter. Requiring $f_{2}$ to reach a certain threshold value produces crossover lines $\lambda^{*}(Q=0)$ that decrease with $\pi_{2}$, as seen in the contour lines on Fig. 5C. This trend holds up to point where consensus is achieved without the need for interdependence, depending on the required threshold value. Beyond this point, the difference in advertising times is enough to impose a majority on the better option, regardless of the particular values of $\pi_{1}$ and $\pi_{2}$.
@@
Figure 6B illustrates the crossover line $\lambda^{*}(\pi_{2};\pi_{1},x=2)$ for different choices of the bad site quality, $q_{1}$. The region where consensus cannot be reached expands as $q_{1}$ increases, pushing the parameters that yield consensus to very high levels of interdependence or a much greater probability to discover the good site. We can determine the value of $\pi_{2}$ from which consensus can be achieved without the need for interdependence. Using the solution at $\lambda=0$ (Eqs. 7 and 8), and the consensus definition, this value is given by:
@@
Finally, we turn our attention to the time required to reach the stationary state, as depicted in Fig. 6D. Once again, we observe the time dilation effect of $\lambda$, coupled with an increase in consensus, indicating a speed-accuracy trade-off. However, in contrast, when $\pi_{1}$ is held constant, increasing $\pi_{2}$ does not exhibit the same effect. Increasing the discovery probability for the good option simplifies the decision problem, achieving better consensus in a shorter time. Interestingly, we notice that in the limit $\pi_{2}\rightarrow 0$, the stationary time increases abruptly around the consensus crossover. This implies that achieving consensus for the best option in such an unfavorable limit comes at the expense of a significantly prolonged decision process. Upon inspecting the characteristic times obtained from the eigenvalues, we observe that this increase in the relaxation time is instead found around the $f_{2}$ transition at $\lambda=\lambda^{\prime}$.
@@ III.3 Finite Size Effects
Figure 7: A: Probability density functions (PDFs) of the stationary site frequencies $f_{1}$ (green) and $f_{2}$ (blue) in simulations with different system sizes $N$. Other parameters are $\pi_{1}=\pi_{2}=0.3$, $q_{1}=7$, $q_{2}=10$ and $\lambda=0.3$. B: Variation of the standard deviation of the PDF for site 2, ($\sigma_{f_{2}}$), across the symmetric discovery parameter space.
@@
Figure 8: Temporal evolution of a single experimental realization in the fully interdependent limit ($\pi_{\alpha}=0\;\forall\alpha$) with equal qualities ($q=10$), starting from different initial conditions. A: $f_{\alpha}(t=0)\simeq(1/3,1/3,1/3)$ B: $f_{\alpha}(t=0)\simeq(0.0,2/5,3/5)$. Other parameters are $\lambda=0.6$, $N=1000$. The black dotted line indicates the value in which the winning population stabilizes, $1-r/\lambda$, while the other vanishes and $f_{0}=r/\lambda$.
The deterministic solution analyzed in the previous subsections aligns perfectly with the results obtained by averaging the stationary state attained in mean-field simulations across all parameter space. However, stochastic simulations enable us to observe significant finite-size fluctuations. An example of such fluctuations for a specific choice of parameters is illustrated in Fig. 7A, where we show the probability density functions (PDFs) of stationary values of the dancing frequencies $f_{1}$ and $f_{2}$ obtained in stochastic simulations on fully connected systems of various sizes ($N$=50,250, and 1000). The variance of the PDFs scales as $N^{-1}$ due to the central limit theorem. As a result, when the system size is relatively small, there is a notable overlap between the PDFs representing the optimal (site 2) and suboptimal (site 1) options. This indicates that a finite stochastic system may temporarily exhibit at least a simple consensus for the suboptimal (bad quality) option ($f_{1}>f_{2}$). However, as the total population increases, the width of these distributions decreases, enhancing the robustness of the decision-making process. T. Galla further characterized these finite-size effects using the van Kampen expansion on the inverse system size [29].
@@
The interplay between independent exploration and social feedback also affects the magnitude of the fluctuations. In the same state space representation as in the previous sections, Fig. 7B illustrates the standard deviation of $f_{2}$ in the symmetric scenario. Excluding some extreme values near or at the fully interdependent limits, the fluctuations are most pronounced at intermediate levels of interdependence. Moreover, these fluctuations appear to increase with $\pi_{1,2}$, akin to the consensus crossover lines discussed earlier.
@@
The main rationale behind this observation lies in the intermediate region of the state space, where the competition between the two primary drivers of the system –independence and interdependence– is most pronounced. High levels of interdependence dominate the system, favoring the optimal site through positive reinforcement facilitated by $\lambda$. Conversely, when $\lambda$ is low, the stationary state is primarily determined by independent discoveries, with minimal or no influence from peer communication. However, in the central region, neither mechanism is dominant enough to outweigh the other, resulting in broader stationary distributions. Despite this broader distribution, consensus may still be established on average.
@@
In the lower-left corner of Fig. 7B, we observe a null variance, followed by a sudden increase at higher values of $\lambda$. This corresponds to the region where the system ends up in a fully uncommitted state, occurring when $\pi_{1,2}=0$ and $\lambda<r_{2}$ in an infinite system. Although analytically, the change in behavior is predicted to occur precisely at $\lambda=r_{2}$, we note that due to finite-size effects, this behavior extends slightly beyond this threshold. At $\lambda\gtrsim r_{2}$, fluctuations suddenly increase as most of the simulation results converge around the predicted stable solution, $f_{0}=r_{2}/\lambda$. However, due to finite-size effects, some results still show the system stabilizing at the fully uncommitted state. As the system size increases, the transition becomes sharper around the threshold $\lambda=r_{2}$.
@@
Lastly, let’s consider the scenario where the site qualities are equal (and their discovery probabilities are also equal) in the fully interdependent limit ($\pi=0$). As we’ve discussed in the analytical solution of the model, detailed in Sec. II.1.2, the system may undergo a phase transition between a fully uncommitted absorbing state and an active state with finite fractions of bees advertising the available sites at $\lambda=r$. We further investigate this non-equilibrium phase transition in the following section. Here, we briefly emphasize the influence of finite-size effects under this limit. While the system may temporarily fluctuate around the stationary state predicted by the deterministic equations, finite-size effects ultimately break this tendency, and the system converges towards a consensus for only one of the options, while the other option disappears. In Figs. 8, we illustrate this phenomenon. In particular, in Figs. 8 A and B, we represent the temporal evolution of $f_{0}$, $f_{1}$ and $f_{2}$ for a binary scenario starting from two different initial conditions. Continuous lines represent the results obtained in stochastic simulations of a finite system consisting of $N=1000$ bees, while dashed lines represent the integration of deterministic equations Eqs. (4), representing the behavior of an infinite system. We observe that fluctuations introduced by finite-size effects ultimately break the symmetry between the two available options. This can occur even when one option initially appears to be favored.
@@ III.4 Absorbing phase transition and finite size effects in the fully interdependent limit
The analytical analysis of the model in Section II.1.3 reveals that in the fully interdependent limit ($\pi_{\alpha}=0$ for all $\alpha$), the system undergoes a phase transition between two states: an absorbing uncommitted state and a partially-active state. In this limit, the system relies entirely on imitation. Thus, if the imitation parameter $\lambda$ exceeds a critical value $\lambda_{c}$, imitation alone is sufficient to sustain a steady state with a finite committed population of bees, denoted as $\rho=\sum_{\alpha\neq 0}f_{\alpha}=1-f_{0}$, at least for the highest quality option. However, below $\lambda_{c}$, no bees actively advertise any site, and the system remains locked in a state where $f_{0}=1$.
@@
Figure 9: Absorbing phase transition in the fully interdependent limit ($\pi_{\alpha}=0,~{}\forall\alpha$) for a fully-connected system. A: Order parameter $\rho=1-f_{0}$ vs. $\tilde{\lambda}$. B: Zoom around the phase transition happening at $\tilde{\lambda}_{c}=1$.
@@
Figure 11: Just noticeable quality difference as a function of the base stimulus strength (A, B) and the system size (C) obtained from deterministic equations. A: Symmetric discovery scenario with $\pi_{1}=\pi_{2}=0.1$, B: Asymmetric discovery scenario with $\pi_{1}=0.15>\pi_{2}=0.05$. C: Simulations for finite system sizes at a fixed $\lambda=0.6$. The discernible quality difference decreases with the system size.
@@ III.5 Weber’s Law of perception
To evaluate whether a swarm accurately discriminates between two stimuli, achieving consensus is necessary. We assess compliance with Weber’s Law by keeping the quality of one option constant while varying the quality of the other until positive consensus is reached. In our experiment, we fixed the high-quality option while adjusting the inferior-quality option, but similar outcomes were obtained when reversing this setup. Figures 11A and B confirm the linear correlation between the difference in quality and the stimulus strength across different combinations of interdependence and discovery probabilities. This analysis was conducted using the original consensus definition, although similar results were obtained with alternative definitions discussed earlier.
@@ III.6 Multiple sites
Figure 12: Consensus parameter $Q$ and dance frequency for the best site $f_{k}$ (shown in the inset) as a function of interdependence for different numbers of options, $k$. The qualities of the sites are defined as follows: $q_{k}=10$, $q_{k-1}=7$, and $q_{k-i}=q_{k-i+1}-1$ for $22\leq i\leq k-$ 2 for options below the top two. A: Each site has a self-discovery probability $\pi_{\alpha}=0.2$. B: Self-discovery probabilities are equal and their sum remains constant, $\sum_{\alpha=1}^{k}\pi_{\alpha}=0.6$, so as the number of sites increases, $\pi_{\alpha}=0.6/k$ decreases.
Up to now, our attention has been directed towards a binary decision problem, where selecting the best option is imperative. Previous research on this model, utilizing different parameter setups, examined a specific scenario involving five options. It’s worth noting that increasing the number of available options doesn’t qualitatively alter the model’s behavior. In a general $k$-site scenario, we still observe the same interplay between independent discovery and social feedback. At high $\lambda$ values, the system effectively identifies the best site, while at low $\lambda$ values, it settles into a multi-opinion state where each site is occupied in proportion to its quality and discovery probability. Therefore, in a broader best-of-N scenario, the specific dance frequency values and consensus transitions will rely on the $\pi_{\alpha}$ and $q_{\alpha}$ parameters defining each unique scenario. However, the overall behavioral trend of the system remains consistent across different setups.
In Fig. 12 we represent the consensus parameter together with the dance frequency for the best-site $f_{k}$ (depicted in the inset) for different number of available sites $k$. We consider two different symmetric scenarios: one where each site has the same self-discovery probability as considered in previous sections, and another where self-discovery probabilities are equal and their sum remains constant, i.e., $\sum_{\alpha=1}^{k}\pi_{\alpha}=0.6$ as in previous sections. In the latter scenario, as the number of sites increases, $\pi_{\alpha}=0.6/k$ decreases. In both scenarios, site qualities are different and ordered following the sequence $q_{1}<q_{2}<...<q_{k}$. While there is a quantitative change with increasing $k$, the system’s behavioral trend observed in the binary problem remains consistent.
@@ A.1 Linear Stability Analysis on the general solution
where $\alpha=1,...k$. The stability of any solution can be verified by obtaining the matrix eigenvalues numerically, once the actual solution $(f_{0}^{,}f_{1}^{,}...,f_{k}^{*})$ is known.
Figure A1: Characteristic relaxation times $\tau$ of linear perturbations obtained from the LSA eigenvalues. A: Symmetric discovery scenario. B: Asymmetric discovery scenario with $\pi_{1}=0.25$. In both scenarios, the qualities are $(q_{1},q_{2})=(7,10)$. White dashed lines indicate the transitions occurring at the fully interdependent limit $\pi_{\alpha}=0$ (null-$\pi_{k}$ limit) for the symmetric (asymmetric) scenario. For the exact value of $\lambda^{\prime}$, refer to Eq. (10).

### v2

@@ III.2 Asymmetric discovery scenario
As before, we can quantify the outcome of the decision process in terms of a consensus parameter. Requiring $f_{2}$ to reach a certain threshold value produces crossover lines $\lambda^{*}$ that decrease with $\pi_{2}$, as seen in the contour lines on Fig. 5C. This trend holds up to point where consensus is achieved without the need for interdependence, depending on the required threshold value. Beyond this point, the difference in advertising times is enough to impose a majority on the better option, regardless of the particular values of $\pi_{1}$ and $\pi_{2}$.
@@
Figure 6B illustrates the crossover line $\lambda^{*}(\pi_{2};\pi_{1},x=2)$ for different choices of the bad site quality, $q_{1}$. The region where consensus cannot be reached expands as $q_{1}$ increases, pushing the parameters that yield consensus to very high levels of interdependence or a much greater probability to discover the good site. We can determine the value of $\pi_{2}$ from which consensus can be achieved without the need for interdependence. Using the solution at $\lambda=0$ (Eqs. (7) and (8)), and the consensus definition, this value is given by:
@@
Finally, we turn our attention to the time required to reach the stationary state, as depicted in Fig. 6D. Once again, we observe the time dilation effect of $\lambda$, coupled with an increase in consensus, indicating a speed-accuracy trade-off. However, in contrast, when $\pi_{1}$ is held constant, increasing $\pi_{2}$ does not exhibit the same effect. Increasing the discovery probability for the good option simplifies the decision problem, achieving better consensus in a shorter time. Interestingly, we notice that in the limit $\pi_{2}\rightarrow 0$, the stationary time increases abruptly around the consensus crossover. This implies that achieving consensus for the best option in such an unfavorable limit comes at the expense of a significantly prolonged decision process. A similar effect is observed when inspecting the characteristic relaxation times obtained from the solutions’ LSA (see Fig. A1). There we observe that this increase in the relaxation time is instead found around the $f_{2}$ transition at $\lambda=\lambda^{\prime}$.
@@ III.3 Finite Size Effects
Figure 7: A: Probability density functions (PDFs) of the stationary site frequencies $f_{1}$ (green) and $f_{2}$ (blue) in simulations with different system sizes $N$. Other parameters are $\pi_{1}=\pi_{2}=0.3$, $q_{1}=7$, $q_{2}=10$ and $\lambda=0.3$. B: Variation of the standard deviation of the PDF for site 2, ($\sigma_{f_{2}}$), across the symmetric discovery parameter space, obtained from simulations with $N=100$ and $q_{1}=7$, $q_{2}=10$.
@@
Figure 8: Temporal evolution of a single experimental realization in the fully interdependent limit ($\pi_{\alpha}=0\;\forall\alpha$) with equal qualities ($q_{\alpha}=10\;\forall\alpha$), starting from different initial conditions. A: $f_{\alpha}(t=0)\simeq(1/3,1/3,1/3)$ B: $f_{\alpha}(t=0)\simeq(0.0,2/5,3/5)$. Other parameters are $\lambda=0.6$, $N=1000$. The black dotted line indicates the value in which the winning population stabilizes, $1-r/\lambda$, while the other vanishes and $f_{0}=r/\lambda$.
[…]
@@
The interplay between independent exploration and social feedback also affects the magnitude of the fluctuations. In the same state space representation as in the previous sections, Fig. 7B illustrates the standard deviation of $f_{2}$ in the symmetric scenario. Excluding some extreme values near or at the fully interdependent limits, the fluctuations are most pronounced at intermediate levels of interdependence. Moreover, the region of high fluctuations appears to shift towards greater $\lambda$ on increasing $\pi_{1,2}$, akin to the consensus crossover lines discussed earlier.
@@
The main rationale behind this observation lies in the intermediate region of the state space, where the competition between the two primary drivers of the system –independence and interdependence– is most pronounced. When high levels of interdependence drive the system, the optimal site is easily favoured by the interplay of longer advertisement and high social feedback. Conversely, when $\lambda$ is low, the stationary state is primarily determined by independent discoveries, with minimal or no influence from peer communication. However, in the central region, neither mechanism is dominant enough to outweigh the other, resulting in broader stationary distributions. Despite this broader distribution, consensus may still be established on average.
@@
In the lower-left corner of Fig. 7B, we observe a null variance, followed by a sudden increase at higher values of $\lambda$. This corresponds to the region where the system ends up in a fully uncommitted state, occurring when $\pi_{1,2}=0$ and $\lambda<r_{2}$ in an infinite system. Although analytically the change in behavior is predicted to occur precisely at $\lambda=r_{2}$, we note that due to finite-size effects, this behavior extends slightly beyond this threshold. At $\lambda\gtrsim r_{2}$, fluctuations suddenly increase as most of the simulation results converge around the predicted stable solution, $f_{0}=r_{2}/\lambda$, while, due to finite-size effects, some results still show the system stabilizing at the fully uncommitted state. As the system size increases, the transition becomes sharper around the threshold $\lambda=r_{2}$.
@@
Lastly, let’s consider the scenario where the site qualities are equal (and their discovery probabilities are also equal) in the fully interdependent limit ($\pi_{\alpha}=0$). As we’ve discussed in the analytical solution of the model, detailed in Sec. II.1.2, the system may undergo a phase transition between a fully uncommitted absorbing state and an active state with finite fractions of bees advertising the available sites at $\lambda=r$. We further investigate this non-equilibrium phase transition in the following section. Here, we briefly emphasize the influence of finite-size effects under this limit. While the system may temporarily fluctuate around the stationary state predicted by the deterministic equations, finite-size effects ultimately break this tendency, and the system converges towards a consensus for only one of the options, while the other option disappears. In Figs. 8, we illustrate this phenomenon, by representing the temporal evolution of $f_{0}$, $f_{1}$ and $f_{2}$ for a binary decision problem starting from two different initial conditions. Continuous lines represent the results obtained in stochastic simulations of a finite system consisting of $N=1000$ bees, while dashed lines represent the integration of deterministic equations Eqs. (4), representing the behavior of an infinite system. We observe that fluctuations introduced by finite-size effects ultimately break the symmetry between the two available options. This can occur even when one option initially appears to be favored.
@@ III.4 Absorbing phase transition and finite size effects in the fully interdependent limit
The analytical analysis of the model in Section II.1.3 reveals that in the fully interdependent limit ($\pi_{\alpha}=0$ for all $\alpha$), the system undergoes a phase transition between two states: an absorbing uncommitted state and a partially-active state. In this limit, the system relies entirely on imitation. Thus, if the imitation parameter $\lambda$ exceeds a critical value $\lambda_{c}$, imitation alone is sufficient to sustain a steady state with a finite committed population of bees, denoted as $\rho=\sum_{\alpha=1}^{\alpha=k}f_{\alpha}=1-f_{0}$, at least for the highest quality option. However, below $\lambda_{c}$, no bees actively advertise any site, and the system remains locked in a state where $f_{0}=1$.
@@
Figure 9: Absorbing phase transition in the fully interdependent limit ($\pi_{\alpha}=0,~{}\forall\alpha$) with equal qualities ($q_{\alpha}=10,~{}\forall\alpha$) for a fully-connected system. A: Order parameter $\rho=1-f_{0}$ vs. $\tilde{\lambda}=\lambda/\lambda_{c}$. B: Zoom around the phase transition happening at $\tilde{\lambda}_{c}=1$.
@@
Refer to captionRefer to caption
Figure 11: Just noticeable quality difference as a function of the base stimulus strength (A, B) and the system size (C). A: Symmetric discovery scenario with $\pi_{1}=\pi_{2}=0.1$, B: Asymmetric discovery scenario with $\pi_{1}=0.15>\pi_{2}=0.05$. Results obtained from analysis of the deterministic equations. C: Simulations for finite system sizes at a fixed $\lambda=0.6$ and $\pi_{1,2}=0.6$. The discernible quality difference decreases with the system size.
@@
Non-equilibrium phase transitions are commonly observed in biological systems [46]. These transitions, which occur far from thermodynamic equilibrium, play a crucial role in shaping the collective behavior of living organisms. Here, we find another example which might be relevant for collective decision making processes.
@@ III.5 Weber’s Law of perception
To evaluate whether a swarm accurately discriminates between two stimuli, achieving consensus is necessary. We assess compliance with Weber’s Law by keeping the quality of one option constant while varying the quality of the other until positive consensus is reached. In our experiment, we fixed the high-quality option while adjusting the inferior-quality option, but similar outcomes were obtained when reversing this setup. Figures 11A and B confirm the linear correlation between the difference in quality and the stimulus strength across different combinations of interdependence and discovery probabilities. This analysis was conducted using the original consensus definition, Eq. (11), although similar results were obtained with alternative definitions discussed earlier.
@@ III.6 Multiple sites
Figure 12: Consensus parameter $Q$ and dance frequency for the best site $f_{k}$ (shown in the inset) as a function of the interdependence for different numbers of options, $k$. The qualities of the sites are defined as follows: $q_{k}=10$, $q_{k-1}=7$, and $q_{k-i}=q_{k-i+1}-1$ for $22\leq i\leq k-$ 2 for options below the top two. A: Each site has a self-discovery probability $\pi_{\alpha}=0.2$. B: Self-discovery probabilities are equal and their sum remains constant, $\sum_{\alpha=1}^{k}\pi_{\alpha}=0.6$, so as the number of sites increases, $\pi_{\alpha}=0.6/k$ decreases.
Up to now, our attention has been directed towards a binary decision problem, where selecting the best option is imperative. Many collective decision making models have been concerned in generalizing their dynamics to multiple option scenarios [37, 24, 39, 40]. Particularly, previous research on this model, utilizing different parameter setups, examined a specific scenario involving five options. It’s worth noting that increasing the number of available options doesn’t qualitatively alter the model’s behavior. In a general $k$-site scenario, we still observe the same interplay between independent discovery and social feedback. At high $\lambda$ values, the system effectively identifies the best site, while at low $\lambda$ values, it settles into a multi-opinion state where each site is occupied in proportion to its quality and discovery probability. Therefore, in a broader best-of-N scenario, the specific dance frequency values and consensus transitions will rely on the $\pi_{\alpha}$ and $q_{\alpha}$ parameters defining each unique scenario. However, the overall behavioral trend of the system remains consistent across different setups.
In Fig. 12 we represent the consensus parameter together with the dance frequency for the best-site $f_{k}$ (depicted in the inset) for different number of available sites $k$. We consider two different symmetric scenarios: one where each site has the same self-discovery probability (independently of the number of sites $k$), and another where the total sum of self-discovery probabilities remains constant, i.e., $\sum_{\alpha=1}^{k}\pi_{\alpha}=0.6$, and consequently each time a new site is added the overall value of $\pi_{\alpha}=0.6/k$ decreases. In both scenarios, site qualities are different and ordered following the sequence $q_{1}<q_{2}<...<q_{k}$. While there is a slight quantitative change with increasing $k$, the system’s behavioral trend observed in the binary problem remains consistent.
@@ A.1 Linear Stability Analysis on the general solution
where $\alpha=1,...k$. The stability of any solution can be verified by obtaining the matrix eigenvalues numerically, once the actual solution $(f_{0}^{*},f_{1}^{*},...,f_{k}^{*})$ is known.
Figure A1: Characteristic relaxation times $\tau$ of linear perturbations obtained from the LSA eigenvalues. A: Symmetric discovery scenario. B: Asymmetric discovery scenario with $\pi_{1}=0.25$. In both scenarios, the qualities are $(q_{1},q_{2})=(7,10)$. White dashed lines indicate the transitions occurring at the fully interdependent limit $\pi_{\alpha}=0$ (null-$\pi_{k}$ limit) for the symmetric (asymmetric) scenario. For the exact value of $\lambda^{\prime}\simeq 0.43$, refer to Eq. (10).